A stationary nucleus of mass 24u emits a ¥ photon. The energy of the emited photon is 7MeV. Find the recoil energy of the nucleus?(ans:-1.1MeV).

Grade:12

1 Answers

vikas askiitian expert
13 years ago

in gama decay there is no change in mass number...

initially nucleous was at rest ,due to internal forces it will move after giving a gamma photon..

so we can apply conservation of momentam here

final momentam = initial momentam = 0

P(gamma photon) + P(nucleous) = 0  .......................1

  energy of photon = 7Mev

   p(photon) = E/c                        (c is speed of light , E =7Mev (given) )

 p(nucleous) = -E/c             ( from eq 1)

  kinetic energy = P2/2m = E2/2mc2 

now put E , C & m ,u will get the  1.1Mev ans